# When Vision Fails: Text Attacks Against ViT and OCR

### v1

@@ 4 Evaluation
TrOCR. TrOCR is a transformer model [26] published by Microsoft implementing OCR [23]. TrOCR achieves state-of-the art performance on text recognition and leverages the modern transformer architecture to implement image understanding and text generation in a single model.
@@
Our first experiment evaluates the performance of TrOCR against diacritics injected via Algorithm 1. For each example, we render the input using the Microsoft Arial Unicode font [22] and pass the rendered image to the TrOCR model for inference. In this experiment, we use negative Levenshtein distance of the model output with the adversarial input as the similarity metric. If the TrOCR model is highly performant, we would expect a trivially small average distance between the input and output when the attack budget is zero (representing no diacritic injections); indeed, this is what we see in the results. The experiments seek to generate adversarial example using diacritics injected into inputs sampled from the grammatically-correct validation split of the CoLA dataset [27], selected as a reproducible sample of short English-language inputs.
@@
In Figure 3 we plot the distance between the output and both the perturbed and unperturbed input. Both measures grow with the budget, indicating that the model neither removes diacritical marks nor recognizes them as diacritics.
@@
In Figure 6 we plot the average chrF score [29] – a measure of translation quality selected as the similarity matric for this experiment – relative to the reference translation as computed by sacreBLEU [30]. The results indicate a decreasing translation quality with increasing diacritics budget.
Defended Toxic Classifier. Similarly, we evaluated another model defended from Unicode attacks via TrOCR. In this experiment, the downstream model is IBM’s toxic content classifier [25]. We use the Wikipedia Detox Dataset [31] as the inputs from which we craft adversarial examples.
The classification performance of the defended model over varying attack budgets is shown in Figure 6. There is, again, a negative relationship between the performance of the model and the attack budget. In fact, with a budget of just two diacritical marks, the performance of the model falls below random (i.e. 50% correct classification).
@@
In Figure 8 we plot the average chrF score – again chosen as the similarity metric – over the perturbation budget. Similar to the TrOCR-defended machine translation experiment, we see a significant negative correlation between translation quality and diacritics perturbation budget.
@@
CANINE Question Answering. Neural encoders are a recent NLP tool designed to replace dictionaries with embeddings learned directly from Unicode code points. Although these models do not unify the visual and encoded pipelines for text input, the learned embeddings claim to offer better flexibility towards otherwise unknown characters. In this experiment, we evaluate whether diacritical mark attacks can subvert models leveraging these neural encoders.
@@
We evaluated a question answering model leveraging the CANINE neural encoder proposed by Google [10]. Specifically, we evaluated a transformer model [33] trained on the SQuAD dataset [34]. Consistent with benchmarks against this dataset, we select F1 score as the similarity metric. The results, shown in Figure 8, show a sharp decline in F1 score with increasing diacritical perturbations.
Figure 4: Performance of evaluated models across perturbation budgets relative to no perturbations.
5 User Study
@@
Having shown that OCR and ViT models perform poorly when interpreting text containing visual perturbations, and that these failures can be used to craft adversarial examples for downstream tasks, we conduct a user survey to understand how human comprehension is affected by diacritical perturbations.
User Study Methodology. The survey contained two main questions, each with multiple sub-components. The questions are included as Figures 10 and 11.
@@
First, to measure human ability to read text containing diacritics, we ask respondents to retype sentences containing diacritics without using any diacritics or other marks. We selected five short, benign sentences as examples to measure comprehension.
@@
Second, to measure human comprehension of text containing diacritics, we asked respondents to identify whether a short sentence is considered toxic using the toxicity definition of Google’s Perspective API [35]. We hand-selected examples to choose short sentences that contained toxic examples with comparatively milder language. We used 12 example sentences in total, evenly split between six toxic and six non-toxic examples. The sentences were randomly ordered within the survey question to avoid biasing respondents.
@@
We surveyed 200 people using Amazon Mechanical Turk, limiting the survey to include only MTurk workers with at least a U.S. high school education since our questions concerned critical interpretation of text. The survey was advertised as “Answer a short (estimated 3-4 min) survey about reading and interpreting short sentences”, intentionally avoiding any mention of diacritics, accents, or unusual marks. We identify six identical low-quality responses that failed quality-control screening, leaving a total of $n=194$ responses.
@@
[Uncaptioned image] True False True 93.8%
[Uncaptioned image] True False True 94.8%
[Uncaptioned image] True False True 96.4%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] False True False 99.0%
[Uncaptioned image] True True True 80.4%
[Uncaptioned image] True True True 97.9%
[Uncaptioned image] True True True 93.3%
@@
[Uncaptioned image] False False False 99.5%
[Uncaptioned image] False False False 99.0%
@@
Table 2: Comparison of OCR and ViT model performance and human performance detecting toxicity in adversarial examples. A ‘True’ label means the adversarial example was identified as toxic, while ‘False’ means the example was considered non-toxic.
We paid respondents \$1.50 USD to complete the survey, which had an average completion time of just under three minutes and maximum completion time of up to ten minutes. This amount was chosen to reflect a fair hourly wage.
@@
User Study Results. We conclude that diacritical marks do not significantly affect a human user’s ability to comprehend the adversarial examples generated in the prior experiments. User responses were highly accurate in both reading and interpreting text containing diacritical marks, indicating that OCR and ViT models struggle with a text interpretation task humans find simple.
@@
We find that the overwhelming majority of human readers are able to correctly read and repeat the example sentences. Respondent accuracy rates for the three example sentences ranged from 93.3% to 97.9% . In evaluating what constitutes an ‘accurate’ response, we ignore minor punctuation differences that seem unlikely to have been caused by diacritical marks (e.g., lack of capitalization, quotation marks around the sentence, a lack of a period at the end of the sentence, etc.).
@@
Inaccurate responses mostly substituted a word that contained a diacritical mark for a similar but incorrect word (e.g., “tube” instead of “tub”, or “John laid” instead of “John lay”). While a small minority of respondents struggled to identify certain words containing diacritics, we conclude that the vast majority of human readers are capable of comprehending text regardless of diacritical marks.
@@
Table 2 contains the twelve examples used in the survey, evenly divided between six examples the TrOCR-defended toxic content classifier labeled correctly (based on the ground-truth toxicity label from the Wikipedia Detox Dataset [31]) and six it did not. Human accuracy was maintained through all adversarial examples suggesting that human accuracy is unaffected by the presence of diacritical marks, even as OCR-defended model accuracy drops.
6 Discussion
Ethics. All attacks performed were run against local models and did not target any online services. We leveraged laboratory GPU resources for attack generation and did not use cloud services. Our user study was reviewed and approved by our Institutional Review Board, and no personal information was tracked from survey participants. We compensated users via Amazon Mechanical Turk at fair market prices. Since the study involved viewing toxic content, per Mechanical Turk policy we included a warning in the survey title that the assignment contains “language that may be considered rude or disrespectful.”
@@ 6 Discussion
Limitations. The adversarial examples in this work produce minor visual artifacts by design. These examples contrast the assumptions made by Boucher et al. [4] in which no visual artifacts were permitted in encoding attacks. However, the results of our user study indicate that the diacritic perturbations produced do not affect the ability of humans reading text, thus giving motivation to break the imperceptibility assumption of encoding attacks.
@@
Defenses. The attacks described in this paper all use Unicode diacritical marks to encode visual perturbations in the textual domain. Diacritics carry important linguistic value – particularly for pronunciation purposes – and cannot be disallowed in inputs without breaking internationalization.
@@
In settings where diacritical marks can’t be removed without hurting the performance of the model, the next best option is to perform adversarial training on the ViT, OCR, or neural encoder processing textual inputs. Yet, this will not come without additional risks via invariance-based adversarial examples where the model will learn to detect characters that are no longer human decipherable [36].
Rendering Design. In the visual text domain, setting-specific engineering details become pseudo-hyperparameters of the model. In this setting, it is no longer sufficient to communicate the model architecture and weights, but rather model providers must provide a thorough implementation or explanation of text rendering, chunking, and canvas handling along with their trained models.
@@
When there is too much text to fit onto the canvas, the canvas cannot be resized to make it fit; this would necessitate the image being further resized before inference, and the text becoming too small for the model to optimally recognize. Instead, long text inputs must be chunked into appropriately sized sections that each fit on their own canvas. These inputs must be processed separately and combined after inference (a process which must account for potential false whitespace caused by the edge of the canvas). Further, if words are broken across canvases the model will likely lose the performance benefits of recognizing adjacent characters of higher likelihood. We provide a visualization of this in Figure 9.
In addition to the aforementioned challenges, the text rendering engine selected must also provide robust rendering functionality for chosen encoding – such as Unicode – inclusive of support for control characters, bidirectional text, and any other details of the text specification.
7 Related Work
Text-Based Adversarial Examples. While most adversarial examples proposed in prior work have been image classification attacks [1, 3, 37], recent work has demonstrated the existence of adversarial examples against text-based ML models. In particular, the accuracy of text classification models decreases on noisy inputs and it is possible to deliberately craft adversarial texts targeting systems reliant on text understanding [11, 12, 13, 14, 38, 39].
In contrast to prior work presenting visibly perceptible perturbations, Boucher et al. [4] previously showed that the discrepancy between text encoding and text visualization can be exploited as an attack vector against text-based machine learning models. They proposed optical character recognition (OCR) models as a “catch-all defense” against such attacks, arguing that the visual nature of the input would mitigate differences in visual and text encoding representation that left text-based processing systems vulnerable.
While Boucher et al. evaluated OCR as a defense against imperceptible perturbations, diacritics present a new class of attacks that are often scarcely more perceptible than invisible perturbations but have a substantial impact on OCR accuracy. Salesky et al. [7] showed that Arabic diacritics and other minor text modifications substantially degrade the performance of text-based models, but concluded that ViTs achieve “relatively robust” performance against diacritized text. In this paper we demonstrate that both OCR systems and ViTs are indeed vulnerable to the Unicode-based adversarial examples introduced by Boucher et al., and that such examples significantly impact the performance of these models.
Attacks on OCR. Most prior work in breaking OCR systems has focused on visual CAPTCHAs [40, 41], but OCR is commonly used in a variety of preprocessing tasks. Recently, Chen et al. [42] demonstrated a class of attacks on OCR using adversarial watermarks in a twist on classic adversarial image examples. The commonality of watermarks on inputs to OCR systems makes this attack effectively imperceptible to humans, though this attack relies on non-text image data in inputs. In contrast, in this work we present a class of attacks in which the adversarial examples are encoded directly into text rather than perturbing an image’s background. To the best of our knowledge, this is the first such attack that manipulates pre-rendered text directly, using the resulting image as an adversarial example.
Toxic Content Detection. Identifying toxic content online has proven to be one of the most pervasive yet elusive research questions in natural language processing. A rapidly growing area of research, toxic content detection is challenging even when classifying standard Unicode characters in the English language, and toxic content classifiers are generally not robust against adversarial examples [43, 44]. Users online commonly deploy a variety of simple text modifications to bypass platform filters: Kurita et al. [43] gives the adversarial example of writing “s*ut up” in the place of ‘shut up’. While most adversarial examples against toxic content detection are visually perceptible, Boucher et al. [4] showed it is also possible to modify the Unicode text encoding of toxic content in a way that is visually imperceptible but fools content detection models. In this paper, we demonstrate that text containing diacritical marks presents a further class of attacks against toxic content detection systems.
8 Conclusion
@@
We have presented a novel method of attacking language models operating in the visual domain by encoding adversarial examples in the textual domain. Once rendered, these examples will generate small visual artifacts that drastically decrease model performance. We accomplish this through the injection of Unicode’s combining diacritical marks via a gradient-free optimization designed to minimize target model performance over the perturbed input. We generate such adversarial examples for real-world models published by Microsoft, Facebook, IBM, and Google finding the attacks to significantly degrade model performance. We also conduct a user study which concludes that the comprehension of human users is not affected by the diacritic injections used in our attacks. In order to defend against such attacks, we conclude that model designers should remove combining diacritical marks from inputs prior to inference.
Appendix
@@
Figure 5: Evaluation of diacritic adversarial examples against FairSeq’s EN->FR translation model defended by TrOCR.
Figure 5: Evaluation of diacritic adversarial examples against FairSeq’s EN->FR translation model defended by TrOCR.
@@
Figure 6: Evaluation of diacritic adversarial examples against IBM’s toxic content detection model defended by TrOCR.
@@
Figure 7: Evaluation of diacritic adversarial examples against the Visual Representation fork of FairSeq for DE->EN translation.
Figure 7: Evaluation of diacritic adversarial examples against the Visual Representation fork of FairSeq for DE->EN translation.
@@
Figure 8: Evaluation of diacritic adversarial examples against SQuAD transformer using Google’s CANINE neural encoder.
Figure 9: An example of three chunking options to render the text “uncharted backwaters of the” onto a series of fixed-size canvases for inference. Option A breaks letters across canvases, which will cause failures during inference. Option B breaks the word backwaters across more canvases than necessary, decreasing the model’s ability to recognize likely adjacencies. Option C is likely the strongest choice in this example, but as with all other options the model pipeline must account for whitespace added by chunking after inference.

### v2

@@ 5.1. TrOCR
TrOCR is a transformer model (Vaswani et al., 2017) published by Microsoft implementing OCR (Li et al., 2021). TrOCR achieves state-of-the art performance on text recognition and leverages the modern transformer architecture.
@@
Our first experiment evaluates the performance of TrOCR against diacritics injected via Algorithm˜1. For each example, we render the input using the Microsoft Arial Unicode font (Microsoft, 2021) and pass the rendered image to the TrOCR model for inference. In this experiment, we use negative Levenshtein distance of the model output with the adversarial input as the similarity metric. If the TrOCR model is highly performant, we would expect - and indeed do see - a small average distance between the input and output when the attack budget is zero (representing no diacritic injections). The experiments seek to generate adversarial examples using diacritics injected into inputs sampled from the grammatically-correct validation split of the CoLA dataset (Warstadt et al., 2018), selected as a sample of short English-language inputs.
@@
In Figure˜5 we plot the distance between the output and both the perturbed and unperturbed input. Both measures grow with the budget, indicating that the model neither removes diacritical marks nor recognizes them as diacritics.
5.2. Defended FairSeq
@@
In Figure˜5 we plot the average chrF score (Popović, 2015) – a measure of translation quality selected as the similarity metric for this experiment – relative to the reference translation as computed by sacreBLEU (Post, 2018). The results indicate a decreasing translation quality with increasing diacritics budget.
Figure 3. Performance of evaluated models across perturbation budgets relative to no perturbations.
5.3. Defended Toxic Classifier
Similarly, we evaluated another model defended from Unicode attacks via TrOCR. In this experiment, the downstream model is IBM’s toxic content classifier (IBM, 2020). We use the Wikipedia Detox Dataset (Thain et al., 2017) as the inputs from which we craft adversarial examples.
The classification performance of the defended model over varying attack budgets is shown in Figure˜7. There is, again, a negative relationship between the performance of the model and the attack budget. In fact, with a budget of just two diacritical marks, the performance of the model falls below random (i.e. 50% correct classification).
5.4. Visual FairSeq
@@
In Figure˜7 we plot the average chrF score over the perturbation budget. Similar to the TrOCR-defended machine translation experiment, we see a significant negative correlation between translation quality and diacritics perturbation budget.
5.5. CANINE Question Answering
@@
Neural encoders are a recent NLP tool to replace dictionaries with embeddings learned directly from Unicode code points. Although these models do not unify the visual and encoded pipelines for text input, the learned embeddings claim to offer better flexibility for unknown characters. In this experiment, we evaluate whether diacritic attacks can subvert models leveraging neural encoders.
@@
We evaluated a question answering model leveraging the CANINE neural encoder proposed by Google (Clark et al., 2022). Specifically, we evaluated a transformer model (Hsu, 2022) trained on the SQuAD dataset (Rajpurkar et al., 2016). Consistent with benchmarks against this dataset, we select F1 score as the similarity metric. The results, shown in Figure˜8, show a sharp decline in F1 score with increasing diacritical perturbations.
6. User Study
@@
Having shown that OCR and ViT models perform poorly when interpreting text containing visual perturbations and that these failures can be used to craft adversarial examples for downstream tasks, we conduct a user survey to understand how diacritical perturbations affect human comprehension.
6.1. User Study Methodology
@@
The survey contained two high-level sections which are each shown in Appendix A.2. First, to measure human ability to read text containing diacritics, we ask respondents to retype sentences containing diacritics, specifying that they should omit diacritics or other marks. We selected 5 short examples to measure comprehension.
@@
Second, to measure human comprehension of text containing diacritics, we asked respondents to identify whether a short sentence is toxic using the toxicity definition of Google’s Perspective API (Google, 2021). For the sentences with a ground-truth toxic label, we filtered to select milder language. We used 12 sentences in total, evenly split between six toxic and six non-toxic examples. The sentences were randomly ordered within the survey question to avoid biasing respondents.
@@
We surveyed 200 people using Amazon Mechanical Turk, limiting the survey to include only workers with at least a U.S. high school education since our questions concerned critical interpretation of text. The survey was advertised as “Answer a short (estimated 3-4 min) survey about reading and interpreting short sentences”, intentionally avoiding any mention of diacritics, accents, or unusual marks. We also warn potential respondents that the survey “contains language that may be considered rude or disrespectful” prior to accepting the task to ensure that they are appropriately informed about the survey contents.
In analyzing the results, we identify six identical low-quality responses that failed quality-control screening, leaving a total of $n=194$ responses. We paid the remaining respondents \$1.50 USD for completing the survey, an amount chosen to reflect a fair hourly wage. The survey had an average completion time of just under three minutes and maximum completion time of up to ten minutes.
6.2. User Study Results
@@
True False True 93.8%
True False True 94.8%
True False True 96.4%
False True False 99.0%
False True False 99.0%
False True False 99.0%
True True True 80.4%
True True True 97.9%
True True True 93.3%
@@
False False False 99.5%
False False False 99.0%
@@
Table 2. Comparison of OCR and ViT model and human performance detecting toxicity in adversarial examples. ‘True’ means the adversarial example was considered toxic, while ‘False’ indicates the example was non-toxic.
@@
Overall, the results demonstrate that diacritical marks do not significantly affect a human user’s ability to comprehend the adversarial examples generated in the prior experiments. User responses were highly accurate in both reading and interpreting text containing diacritical marks, indicating that OCR and ViT models struggle with a text interpretation task humans find simple.
6.2.1. Text Legibility
@@
We find that the overwhelming majority of human readers are able to correctly read and repeat the example sentences, with respondent accuracy rates ranging from 93.3% to 97.9%. In evaluating what constitutes an ‘accurate’ response, we ignore minor punctuation differences unlikely to have been caused by diacritical marks (e.g., lack of capitalization, quotation marks around the sentence, a lack of a period at the end of the sentence).
@@
Inaccurate responses mostly substituted a word that contained a diacritical mark for a similar but incorrect word (e.g., “tube” instead of “tub”, or “John laid” instead of “John lay”). In such cases, we take a conservative approach and err on the side of marking it as incorrect since we cannot determine whether the substitution was due to the presence of diacritical marks or to ordinary typographical human errors. We conclude that while a small minority of respondents struggled to identify certain words containing diacritics, the vast majority of human readers are readily capable of comprehending text regardless of diacritical marks.
6.2.2. Text Comprehension
@@
In the second question, we test individuals’ ability to identify toxic text as a proxy for human comprehension of the text. Table˜2 contains the 12 examples used in the survey, evenly divided between six examples the TrOCR toxic content classifier labeled correctly (based on the ground-truth toxicity label from the Wikipedia Detox Dataset) and six it did not. Human readers were able to correctly classify all 12 sentences, including the six examples that the TrOCR model classified incorrectly, with respondent accuracy ranging from 80.4% to 99.0% for a given sentence.
In some cases, the percentage of human readers identifying a sentence as toxic was seemingly affected more by individual perceptions of toxicity than by the diacritical marks. For instance, the adversarial example “this band sucks anyways” was correctly identified as toxic by only 80.4% of respondents, a noticeably lower consensus than the other examples (the sentence with the second-lowest performance had an accuracy of 93.3%) though still comprising a clear majority. Given the high consensus for all other examples, it seems likely this discrepancy is largely due to users viewing the statement as comparatively mild and therefore perceiving it as non-toxic. Several sentences that were unambiguously non-toxic, such as “John often meets Mary”, saw a human accuracy rate of 99.0%, supporting this hypothesis.
Given that human accuracy was easily maintained through all examples, we conclude that human performance is unaffected by the presence of diacritical marks even as OCR-defended model performance drops, demonstrating that OCR models perform far worse than baseline human comprehension. While our primary attack scenarios assume automated input processing where a human might not look at the text inputs at any point, given the effectiveness of the attack even with just one perturbation as shown in Figure˜3 the attack could easily go unnoticed or unremarked even if the text was observed. It would be entirely feasible for an adversary to carefully select examples with minimal perturbation that are nearly indistinguishable from the original text.
7. Discussion
7.1. Limitations
@@
The adversarial examples in this work produce minor visual artifacts by design. These examples contrast the assumptions made by Boucher et al. (2022) in which no visual artifacts were permitted in encoding attacks. However, the results of our user study indicate that the diacritic perturbations produced do not affect the ability of humans reading text, thus giving motivation to break the imperceptibility assumption of encoding attacks.
7.2. Defenses
@@
The attacks described in this paper all use Unicode diacritical marks to encode visual perturbations in the textual domain. Diacritics carry important linguistic value – particularly for pronunciation purposes – and cannot be disallowed in inputs without breaking internationalization.
@@
In settings where diacritical marks can’t be removed without hurting the performance of the model, the next best option is to perform adversarial training on the ViT, OCR, or neural encoder processing textual inputs. Yet, this will not come without additional risks via invariance-based adversarial examples where the model will learn to detect characters that are no longer human decipherable (Tramèr et al., 2020).
7.3. Rendering Design
In the visual text domain, setting-specific engineering details become pseudo-hyperparameters of the model. In this setting, it is no longer sufficient to communicate the model architecture and weights, but rather model providers must provide a thorough implementation or explanation of text rendering, chunking, and canvas handling along with their trained models.
@@ 7.3. Rendering Design
When there is too much text to fit onto the canvas, the canvas cannot be resized to make it fit; this would necessitate the image being further resized before inference, and the text becoming too small for the model to optimally recognize. Instead, long text inputs must be chunked into appropriately sized sections that each fit on their own canvas. These inputs must be processed separately and combined after inference (a process which must account for potential false whitespace caused by the edge of the canvas). Further, if words are broken across canvases the model will likely lose the performance benefits of recognizing adjacent characters of higher likelihood. We provide a visualization of this in Appendix A.3.
8. Ethics
All attacks were run against local models and did not target any online services. We leveraged laboratory GPU resources for attack generation and did not use cloud services. The attack itself, however, could be applied to any external service taking in text-based user input. We notified Google of our findings on 18.05.2023 but they opted not to make changes to their transformer model, CANINE, at the time. We notified Microsoft, Meta, and IBM around the same time but have not heard back as of this writing.
Our user study was approved by our Institutional Review Board, and no personal information was tracked from survey participants. We compensated users via Amazon Mechanical Turk at fair market prices. Since the user study involved viewing toxic content, we included a warning in the survey title that the assignment contains “language that may be considered rude or disrespectful.”
9. Conclusion
@@
We present a novel method of attacking language models operating in the visual domain by encoding adversarial examples in the textual domain. Once rendered, these examples will generate small visual artifacts that drastically decrease model performance. We accomplish this through the injection of Unicode’s combining diacritical marks via a gradient-free optimization designed to minimize target model performance over the perturbed input. We generate such adversarial examples for real-world models published by Microsoft, Facebook, IBM, and Google finding the attacks to significantly degrade model performance. We also conduct a user study which concludes that the comprehension of human users is not affected by the diacritic injections used in our attacks. In order to defend against such attacks, we conclude that model designers should remove combining diacritical marks from inputs prior to inference.
Appendix A Appendix
A.1. Metric-Specific Experimental Results
Figure 4. Evaluation of diacritic adversarial examples against TrOCR.
@@ A.1. Metric-Specific Experimental Results
Figure 5. Eval of diacritic adv. examples against FairSeq EN->FR translation defended by TrOCR.
@@
Figure 6. Eval of diacritic adv. examples against IBM’s toxic content detection model defended by TrOCR.
@@
Figure 7. Eval of diacritic adv. examples against Visual Representation fork of FairSeq for DE->EN translation.
@@
Figure 8. Eval of diacritic adv. examples against SQuAD transformer using Google’s CANINE.
A.2. User Study Questions